# Bidimensional Exploration of the warm-Temperature Ionised gaS (BETIS)

### v1

@@ Appendix D ESO443-21
![Refer to caption](extracted/5688766/ESO443-21_eDIG_map_SII_NII.jpg)
![Refer to caption](extracted/5688766/ESO443-21_eDIG_map_OIII_OI.jpg)
@@
![Refer to caption](extracted/5688766/lines_z_ESO443-21.jpg)
![Refer to caption](extracted/5688766/lines_GCD_up_ESO443-21.jpg)
@@
![Refer to caption](extracted/5688766/ESO443-21_S_T_GCD.jpg)
![Refer to caption](extracted/5688766/ESO443-21_SII_OI_GCD.jpg)
@@
![Refer to caption](extracted/5688766/ESO443-21shocks_Z0.04_q1e7_0.8_-1_2.jpg)
![Refer to caption](extracted/5688766/ESO443-21shocks_Z0.04_q1e7_0.7_down.jpg)
@@ Appendix E ESO469-15
![Refer to caption](extracted/5688766/ESO469-15_eDIG_map_SII_NII.jpg)
![Refer to caption](extracted/5688766/ESO469-15_eDIG_map_OIII_OI.jpg)
![Refer to caption](extracted/5688766/lines_z_ESO469-15.jpg)
The galaxy ESO469-15 exhibits a noticeable presence of discrete extraplanar H II regions (also reported by other authors, e.g. Rautio et al. 2022). Before performing the analysis, we masked all the extraplanar H II regions using the python routine DAOStarFinder from the PhotUtils555DAOStarFinder documentation. package. Figure 35 shows the line maps after the H II regions removal. The line ratios near to the H II regions are even lower than those close to the bright galactic centre, being considerably low compared to the rest of the eDIG, specially at z ¡ 0, where most of the extraplanar H II regions are located. The effect of the extraplanar H II regions is clear in the line distributions (Figure 36). The trend of the integrated distribution changes continuously along the distance from the midplane, and squares distributions (at MAD = 1.5, where most of the extraplanar H II regions are located) present considerably lower line ratios in comparison with the rest of the distributions at z ¡ 0, up to 0.6 for the [S II] /H and [N II] /H ratios. [O III] /H ratio shows a mixed tendency for every distribution, as for the rest of the galaxies. Finally, the [O I] /H ratio shows lower values at MAD = 0 kpc and z 0 kpc, since for at higher distances the S/N of the [O I] line is not sufficient to sample those regions.
![Refer to caption](extracted/5688766/lines_GCD_up_ESO469-15.jpg)
@@
![Refer to caption](extracted/5688766/ESO469-15shocks_Z0.04_q1e7_0.7.jpg)
![Refer to caption](extracted/5688766/ESO469-15_S_T_GCD.jpg)
![Refer to caption](extracted/5688766/ESO469-15_SII_OI_GCD.jpg)
@@ Appendix F ESO157-49
![Refer to caption](extracted/5688766/ESO157-49_eDIG_map_SII_NII.jpg)
![Refer to caption](extracted/5688766/ESO157-49_eDIG_map_OIII_OI.jpg)
@@
![Refer to caption](extracted/5688766/lines_z_ESO157-49.jpg)
![Refer to caption](extracted/5688766/lines_GCD_up_ESO157-49.jpg)
@@
![Refer to caption](extracted/5688766/ESO157-49shocks_Z0.02_q1e7_0.5.jpg)
![Refer to caption](extracted/5688766/ESO157-49shocks_Z0.04_q1e7_0.5_0.5.jpg)
![Refer to caption](extracted/5688766/ESO157-49shocks_Z0.04_q1e7_0.7_2.5.jpg)
![Refer to caption](extracted/5688766/ESO157-49_S_T_GCD.jpg)
![Refer to caption](extracted/5688766/ESO157-49_SII_OI_GCD.jpg)
@@ Appendix G IC217
![Refer to caption](extracted/5688766/IC217_eDIG_map_SII_NII.jpg)
![Refer to caption](extracted/5688766/IC217_eDIG_map_OIII_OI.jpg)
![Refer to caption](extracted/5688766/lines_z_IC217.jpg)

### v2

@@ Appendix E ESO469-15
The galaxy ESO469-15 exhibits a noticeable presence of discrete extraplanar H II regions (also reported by other authors, e.g. Rautio et al. 2022). Before performing the analysis, we masked all the extraplanar H II regions using the python routine DAOStarFinder from the PhotUtils777DAOStarFinder documentation. package. Figure 35 shows the line maps after the H II regions removal. The line ratios near to the H II regions are even lower than those close to the bright galactic centre, being considerably low compared to the rest of the eDIG, specially at z ¡ 0, where most of the extraplanar H II regions are located. The effect of the extraplanar H II regions is clear in the line distributions (Figure 36). The trend of the integrated distribution changes continuously along the distance from the midplane, and squares distributions (at MAD = 1.5, where most of the extraplanar H II regions are located) present considerably lower line ratios in comparison with the rest of the distributions at z ¡ 0, up to 0.6 for the [S II] /H and [N II] /H ratios. [O III] /H ratio shows a mixed tendency for every distribution, as for the rest of the galaxies. Finally, the [O I] /H ratio shows lower values at MAD = 0 kpc and z 0 kpc, since for at higher distances the S/N of the [O I] line is not sufficient to sample those regions.